Help me solve this

Exercise 18-11 Your answer is incorrect. Try again. Here is the income statement for Myers, Inc. Myers, Inc. Income Statement For the Year Ended December 31, 2020 Sales revenue Cost of goods sold Gross profit Expenses (including $10,000 interest and $30,000 income taxes) Net income $404,000 186,000 218,000 98,000 $ 120,000 Additional information: 1. 2. 3. Common stock outstanding January 1, 2020, was 16,000 shares, and 24,000 shares were outstanding at December 31, 2020. The market price of Myers stock was $11.20 in 2020. Cash dividends of $25,200 were paid, $8,000 of which were to preferred stockholders.

Compute the following measures for 2020. (Round Earnings per share to 2 decimal places, e.g. 1.65, and all other answers to 1 decimal place, e.g. 6.8 or 6.8%.) (a) Earnings per share 11.20 (b) Price-earnings ratio (c) Payout ratio (d) Times interest earned Click if you would like to Show Work for this question: times times Open Show Work
Answer #1

All workings forms part of the answer

Requirement 'a'

A Net Income $              120,000
B Preferred stock dividend $                   8,000
C = A - B Net Income for Common Stockholder $              112,000
D Beginning common stock outstanding                     16,000
E Ending common stock outstanding                     24,000
F = (D+E)/2 Average common stock outstanding                     20,000
G = C/F Earnings per share $                     5.60

* Requirement 'b'

A Market price per share $                   11.20
B Earnings per share $                     5.60
C = A/B Price earnings ratio                            2.0

* Requirement 'c'

A Total dividend to common stockholders $                17,200
B Average common stock outstanding                     20,000
C= A/B Dividend per share $                     0.86
D Earnings per share $                     5.60
E = C/D Payout ratio 15.4%

* Requirement 'd'

A Net Income $              120,000
B Interest expense $                10,000
C Income taxes $                30,000
D = A+B+C Earnings before interest and tax $              160,000
E = D/B Times interest earned                         16.0